A car heading east through a city accelerates after passing a signpost marking the city limits. It is accelerating at


constant magnitude of 3.0 m/s2. At time t = 0 the car is 5.0 m east of the signpost, moving east at 15 m/s. a) Find the car’s position at t=2.0 s.


b) Find the car’s velocity at t=2.0 s.


c) Where is the car when its velocity is 25 m/s.


1
Expert's answer
2022-03-02T10:07:15-0500

a)

"s=s_0+v_0t+\frac{at^2}2=41~m,"

b)

"v=v_0+at=21~\frac ms,"

c)

"s_1=s_0+\frac{v^2-v_0^2}{2a}=71.7~m."